A= C + + G + X - M C = 0.75Y = 1200 - 50 G = 100 X = 300 M= 200 Y = A L = 0.25Y - 25i (M/P) = 250 L = (M/P) Derive the IS equI need help with the last part of this questions with the new curve etc
Answer #1

If Govt Increases its spending by 100, then New is will be a 100 too new a = 200 New Is equation y=C+I+ 4+x-M Y 20.757 +1200S000 = 1201 is sooo 120 ů = 41.66 Substituting i into Is equation y = 6000 - 20 (U1.66) = 6000- 83333 Y = 5,166.66 iv) - Aty